OpenStudy (anonymous):

If the probability density of a random variable is given by: f(x)= K(x+1) for -1

OpenStudy (amistre64):

K(x+1) is just some function that you have to prolly solve for K to define it as a probability

OpenStudy (amistre64):

you integrate f(x) as it is defined, and integrate it along the stated interval .. K is just a constant. Solve for K so that the integration is equal to 1

OpenStudy (anonymous):

I keep getting the wrong answer when I solve it, my key says it should be 1/2 but I am getting 1 as my answer

OpenStudy (anonymous):

@amistre64 ^^^^^

OpenStudy (amistre64):

int K(x+1) from -1 to 1 int Kx + K from -1 to 1 Kx^2/2 + Kx from -1 to 1 = 1 K/2 + K - K/2 + K = 1 K(1/2 + 1 - 1/2 + 1) = 1 K(2) = 1 K = 1/2

OpenStudy (amistre64):

\[\int_{1}^{-1}K(x+1)~dx=1\] solve for the constant K

OpenStudy (anonymous):

This is how I set it up \[\int\limits_{-1}^{1}(\frac{ x^2 }{ 2 }-1)\]

OpenStudy (amistre64):

why would you setup up something different then what the information in the question provides?

OpenStudy (anonymous):

idk that how our teacher taught us to solve it. he told us to set it up so that k is on the outside of the \[\int\limits_{-1}^{1}\]

OpenStudy (amistre64):

then how does K on the outside change (x+1) into (x^2/2 - 1) ?

OpenStudy (amistre64):

\[\int_{1}^{-1}K(x+1)~dx=1\] \[K\int_{1}^{-1}x+1~dx=1\]

OpenStudy (anonymous):

and then you integrate the (x+1), im confused on how to integrate the +1. like if the equation was \[\int\limits_{-1}^{1}kx(x+1)\] then you would get \[\int\limits_{-1}^{1}(x^2)+x\]

OpenStudy (anonymous):

the integrate will give you \[\int\limits_{-1}^{1}(\frac{ x^3 }{ 3 }-\frac{ x^2 }{ 2 })\] idk thats just what i have in my notes i was trying to follow and then you solve for k

OpenStudy (anonymous):

hmm i think I see what your talking about I typed a negative sign by mistake???

OpenStudy (amistre64):

the probability density of a random variable is given by: f(x)= K(x+1) for -1<x<1 so your notes have the modified version you are presenting? If so, work it on the information in the stated problem. Is it that you dont know how to integrate? or that you are simply integrating the wrong thing?

OpenStudy (anonymous):

i am not sure with one of the rules of integration when dealing with a whole number. I know that if I have a variable I plus one to that and divide by that plus one. but with a whole number do I do that also or just leave the number?

OpenStudy (amistre64):

the constant rule of integration eh .. consider a derivative that produces a constant: nx, for some constant n derives to: n if n=1 thats just: 1x derives to 1 if n=4 thats just: 4x derives to 4 etc

OpenStudy (amistre64):

you could just do a clean up .. let u = (x+1) , then du = dx

OpenStudy (amistre64):

\[\int Ku~du\] \[K\int u~du\] \[K(\frac12 u^2)~:~[a,b]\]

OpenStudy (amistre64):

im getting the thought that you are taking a stats class, and are rusty or have not taken a calc 2 class maybe?

OpenStudy (anonymous):

yea I haven't taken calc in 3 years so i am rusty and in a stats class..lol

OpenStudy (amistre64):

:) the +1 just integrates up to x is all if we want to go that route.

OpenStudy (anonymous):

Thank you soooooo much!!!!!

OpenStudy (amistre64):

good luck :)